OpenStudy (anonymous):

If sin(x) = A, find the value of sin(2π - x). My answers are A, -A, 2A and 1/2A. I think it's 2A, am I correct? Need help asap.

OpenStudy (anonymous):

@ganeshie8 would you mind checking this for me? :)

ganeshie8 (ganeshie8):

sin(2pi -x) = sin(-x) = -sin(x) = -A

OpenStudy (anonymous):

and you are correct yet again, smarty pants :) thank you so much @ganeshie8

OpenStudy (anonymous):

If cos(x) = A, then cos (x + π)=

ganeshie8 (ganeshie8):

lets use this formula : cos(A+B) = cosA cosB - sinA sin B

ganeshie8 (ganeshie8):

cos(x + pi) = cos(x) cos(pi) - sin(x) sin(pi)

ganeshie8 (ganeshie8):

see if u can simplify :)

OpenStudy (anonymous):

it's -A again, right?

ganeshie8 (ganeshie8):

Yes !!

OpenStudy (anonymous):

yay!!! last one? If tan(x) = N, then tan(x + π) =

ganeshie8 (ganeshie8):

il give a hint : period of tanx = pi

ganeshie8 (ganeshie8):

the value of tan repeats every pi, so, tan(x + pi) = tan(x) = N

ganeshie8 (ganeshie8):

if that makes sense.. . :)

OpenStudy (anonymous):

So would it just be N?

ganeshie8 (ganeshie8):

yes

OpenStudy (anonymous):

yay!!! thank you sooo much, always appreciated :)

ganeshie8 (ganeshie8):

np.. you wlcme :)